Hakim invested $15,000. He put part in a term deposit that paid 4% per annum, and the remainder in T - bills at 5% per annum. After one year the total interest earned was $690. How much did Hakim invest at each rate?
Answer by Boreal(15235)

x=4% investment amount
15000-x= 5% investment amount
690=.04x + .05 (15000-x) since the interest earned is the principal *rate* time, which is one year, or 1
690=0.04x+ 750-.05x
-60=0.01x
multiply by (-100)
$6000=x That makes $240 at 4%
$9000=15000-x. That makes $450 at 5% ; total interest is 690. Checks.